OpenStudy (anonymous):

In the below system, solve for y in the first equation. x + 3y = 6 2x – y = 10

jimthompson5910 (jim_thompson5910):

Subtract x from both sides, then divide both sides by 3 to isolate y \[\Large x + 3y = 6\] \[\Large 3y = 6-x\] \[\Large y = \frac{6-x}{3}\]

jimthompson5910 (jim_thompson5910):

If you wanted to solve this system an easier way, I'd solve for y in the second equation (to avoid fractions)

OpenStudy (anonymous):

my choices are 1/3x+2 -1/3x+6 -x+2 -1/3x+2 ?

jimthompson5910 (jim_thompson5910):

you can then simplify the last equation \[\Large y = \frac{6-x}{3}\] \[\Large y = \frac{6}{3}-\frac{x}{3}\] \[\Large y = 2-\frac{1}{3}x\] \[\Large y = -\frac{1}{3}x+2\] So the answer is choice D

OpenStudy (anonymous):

thnks so much